OpenStudy (anonymous):

A variable force of 4x−2 pounds moves an object along a straight line when it is x feet from the origin. Calculate the work done in moving the object from x = 1 ft to x = 18 ft. (Round your answer to two decimal places & a medal will be given for good steps as always :-) )

OpenStudy (anonymous):

Ack that's supposed to be the force being: \[4x^{-2}\] Sorry for any confusion this typo caused

terenzreignz (terenzreignz):

\[\int\limits_{1}^{18}4x ^{-2}dx\]

OpenStudy (anonymous):

That's it? Hmm... I thought it would have been a far more complicated setup

terenzreignz (terenzreignz):

Actually, that's as far as I can go (besides evaluating the integral, of course) I honestly can't explain why this is the set up T.T

OpenStudy (anonymous):

I just keep getting the integral does not converge :-/

terenzreignz (terenzreignz):

Besides, if the force wasn't variable, ie, a constant c, moved from a to b, then \[\int\limits_{a}^{b}cdx\] would still give the work done :)

OpenStudy (anonymous):

And yes, I'm certain that's what it says: \[4x^{-2} lbs\]

terenzreignz (terenzreignz):

pound is a unit of force, right?

OpenStudy (anonymous):

Indeed it is. Same as a newton in terms of what it's measuring (even though they are different values)

OpenStudy (anonymous):

I would prefer it was in metric tbh lol

terenzreignz (terenzreignz):

So, was I right? do I get the remuneration? ;)

OpenStudy (anonymous):

\[lbs \neq mass\]

OpenStudy (anonymous):

Well... It's not convergent though

OpenStudy (anonymous):

So how the hell am I supposed to get a quantified answer if it doesn't converge, very odd

OpenStudy (anonymous):

But I keep re-reading the question and that is exactly, word for word, what it says

OpenStudy (anonymous):

No instructions are forthcoming for a lack of covergence

OpenStudy (anonymous):

\[\int\limits 4/x^2 dx = -4/x+constant\]

OpenStudy (anonymous):

I'll be back to check in on this tomorrow :-)

terenzreignz (terenzreignz):

\[\int\limits_{1}^{18}4x ^{-2}dx\] \[\int\limits_{}^{}4x ^{-2}dx = -4x ^{-1} + C\] -4/x, evaluated from 1 to 18 -4/18 - (-4/1) = 4 - 2/9 = 34/9

terenzreignz (terenzreignz):

I don't really see what's wrong, am I missing something here?

terenzreignz (terenzreignz):

I think the answer is 34/9

OpenStudy (anonymous):

You are correct and ty for showing steps so I can follow along! :-D

terenzreignz (terenzreignz):

Haha, thanks, and no problem :)

OpenStudy (anonymous):

Nah, you're good, this is the easy part: \[ \int\limits_{}^{}4x ^{-2}dx = -4x ^{-1} + C \]

OpenStudy (anonymous):

-2+1 = -1